OpenStudy (anonymous):

Let F(s)=5s^2+5s+4. Find a value of d greater than 0 such that the average rate of change of F(s) from 0 to d equals instantaneous rate of change of F(s) at s=1.

OpenStudy (anonymous):

this is a similar prob to what i posted earlier but im getting the wrong answer so im doing something wrong

OpenStudy (jamesj):

The average rate of change of F(s) from 0 to d is the average of F'(s) from 0 to d, which is \[ \frac{1}{d} \int_0^d dF/ds \ ds = \frac{F(d) - F(0)}{d} \] Hence you want to find a d such that \[ \frac{F(d) - F(0)}{d} = F'(1) \] So now you just need to evaluate.

OpenStudy (jamesj):

talk to me

OpenStudy (anonymous):

hold on. im processing this information. lol

OpenStudy (anonymous):

so 5d+5=-2/5 ???

OpenStudy (jamesj):

5d + 5 I agree with. Calculate F'(1) again.

OpenStudy (anonymous):

idk what im doing wrong

OpenStudy (anonymous):

10s+5=1? no?

OpenStudy (jamesj):

No, you want the value of dF/ds when s = 1. That's what F'(1) means.

OpenStudy (jamesj):

F'(s) = 10s + 5. Hence F'(1) = ...?

OpenStudy (anonymous):

15

OpenStudy (jamesj):

yes

OpenStudy (jamesj):

Thus 5d + 5 = 15

OpenStudy (anonymous):

i wasnt plugging in i was setting equal to 1

OpenStudy (anonymous):

so the answer is 2

OpenStudy (jamesj):

Would seem so!

OpenStudy (anonymous):

yeah, it is